# Louis Pallegoix1, Jaime Travesedo1, Alexandre S. May1,2, Léo Balembois1, Denis Vion1, Patrice Bertet*1 and Emmanuel Flurin Université Paris-Saclay, CEA, CNRS, SPEC, 91191 Gif-sur-Yvette Cedex, France 2Alice&\&&Bob, 53 boulevard du Général Martial Valin, 75015 Paris

### v1

@@ Acknowledgements
After the etching step, the chip shall be re-cleaned before moving forward to the e-beam related steps. Recently, we have added an oxide removal step in the process which has not been applied to the sample presented here. First of all, the sample is dipped in a bath of acetone during 5 minutes for removing AZ1518, then in another bath of IPA during 3 minutes, then finally in piranha solution during 2 minutes. The chip is rinsed with water, dehydrated with IPA and dried by dinitrogen blow.
In the recently added oxidation step, the chip is dipped during 20 minutes in buffered oxide etchant (BOE), rinsed with a first bath of water during 20 seconds with agitation, then in a second bath of water for transport between benches, dehydrated with IPA, dried by dinitrogen blow and dehydrated further on hotplate during 5 minutes at 120°C.
@@
A first layer made of MAA EL-6 is spincoated during 60 seconds at 4000 rpm with an acceleration of 1000 rpm/s. Postbake is performed on hotplate during 5 minutes at 172°C. Then comes the spincoating of a second layer made of PMMA 950k A6 during 60 seconds at 3000 rpm with an acceleration of 2000 rpm/s. Postbake, on a hotplate at 175°C, lasts 15 minutes.
The substrate being made of sapphire, there is no way the electrons escape through the substrate during lithography. Therefore, the bilayer shall be metallised for preventing charge effects. In that regard, a layer of 7 nm of aluminum is evaporated over the bilayer at a rate of 0.2 nm/s. There is no need for a good vacuum for this evaporation step, a pressure of about $10^{-6}$ hPa is enough.
@@
Given the need of the simultaneous presence of Josephson junctions of very different sizes over the device, the Manhattan process has been chosen. Due to the absence of a 100 kV EBL in our laboratory, we have developed an adapted version of this process using a bilayer with a reversed aspect ratio (a layer of MAA about 170 nm thick under a layer of PMMA about 580 nm thick). With our Raith e-LINE system, we are using the following parameters :
  • Magnification : ×$\times 1000$
  • Base dose : $250\mu C/cm^{2}$
@@
  • Voltage : 25 kV
  • Aperture : 10 µm
  • Current : About 20 pA
  • Writefield : $100\times 100\mu m^{2}$
C6 : Aluminum etching and resist development and descum
Before developing the e-beam resist, one shall remove the metallisation layer. Aluminum etching is performed by dipping the chip in a solution of potassium hydroxyde (KOH) of 10 g/L in water during 30 seconds. It is then rinsed by dipping 3 seconds in water, then 3 seconds in IPA. The development itself begins right after : The chip is dipped during 45 seconds in a solution made of 1/4 methylisobutylketone (MIBK) and 3/4 IPA (in volume). Development is stopped by dipping in IPA during 15 seconds. The chip and the bilayer are dried by nitrogen blow first, then on a hotplate at 70°C during 1 minute. The resist mask is finally cleaned in a dioxygen plasma asher during 10 seconds, with a pressure of 0.2 hPa and an ionising power of 75 W.
@@
For evaporating Josephson junction, a good vacuum is required (at least below $5\cdot 10^{-7}$ hPa). Given that we are using the Manhattan process, the evaporator shall be featured with a planetary rotation. The parameters informed here work for one of our Plassys MEB 550S system.
First of all, the surface is prepared by argon milling with an incidence angle of 0°. The ion gun is tuned with filament submitted to 500 V and 25 mA, and acceleration grid to 100 V. Shutter is off ion beam trajectory 2 times 5 seconds, with 1 minute of rest between each. Then comes the evaporation of the first layer of aluminum. It is performed with an angle of 30° and at a rate of 1 nm/s. A thickness of 29 nm is asked to the evaporator in order to obtain an actual thickness of 25 nm. This first layer is oxidised by a mixture of dioxygen and argon at 10 hPa during 5 minutes. The sample holder is rotated around the planetary axis of 90° in order to evaporate along the orthogonal direction of the Manhattan pattern. Finally, the second evaporation of aluminum is performed with the same parameters (30° angle and 1 nm/s rate) except that a thickness of 69 nm is asked for in order to get an actual thickness of 60 nm.

### v2

@@ Acknowledgements
After the etching step, the chip is re-cleaned before proceeding to the e-beam related steps. Recently, we have added an oxide-removal step to the process which has not been applied to the sample presented here. First, the sample is immersed in a bath of acetone for 5 minutes to remove AZ1518, then in another bath of IPA for 3 minutes, and finally in piranha solution for 2 minutes. The chip is rinsed with water, dehydrated with IPA and dried with a stream of nitrogen.
In the recently added oxidation step, the chip is immersed for 20 minutes in buffered oxide etchant (BOE), rinsed in a first bath of water for 20 seconds with agitation, then in a second bath of water for transport between benches, dehydrated with IPA, dried with a stream of nitrogen, and further dehydrated on a hotplate during 5 minutes at 120°C.
@@
A first layer of MAA EL-6 is spin-coated for 60 seconds at 4000 rpm with an acceleration of 1000 rpm/s. The postbake is performed on a hotplate for 5 minutes at 172°C. This is followed by spin coating of a second layer of PMMA 950k A6 for 60 seconds at 3000 rpm with an acceleration of 2000 rpm/s. The postbake, on a hotplate at 175°C, lasts 15 minutes.
The substrate being made of sapphire, there is no way for electrons to escape through it during lithography. Therefore, the bilayer must be metallized to prevent charge effects. For this reason, a 7-nm thick layer of aluminum is evaporated over the bilayer at a rate of 0.2 nm/s. There is no need for a good vacuum for this evaporation step ; a pressure of about $10^{-6}$ hPa is enough.
@@
Given the need for the simultaneous presence of Josephson junctions of very different sizes across the device, the Manhattan process has been chosen. Due to the absence of a 100-kV electron beam lithography (EBL) system in our laboratory, we developed an adapted version of this process using a bilayer with a reversed aspect ratio (a layer of MAA about 170 nm thick beneath a layer of PMMA about 580 nm thick). We use the following parameters with our Raith e-LINE system :
  • Magnification : $\times 1000$.
  • Base dose : $250\mu C/cm^{2}$.
@@
  • Voltage : 25 kV.
  • Aperture : 10 µm.
  • Current : About 20 pA.
  • Writefield : $100\times 100\mu m^{2}$.
C6 : Aluminum etching, resist development, and descum
Before developing the e-beam resist, the metallization layer must be removed. Aluminum etching is performed by dipping the chip in a 10 g/l potassium hydroxyde (KOH) solution in water for 30 seconds. It is then rinsed by for 3 seconds in water, followed by 3 seconds in IPA. The development itself begins immediately after : The chip is immersed for 45 seconds in a solution made of 1/4 methyl isobutyl ketone (MIBK) and 3/4 IPA (by volume). Development is stopped by dipping in IPA for 15 seconds. The chip and the bilayer are dried first with a stream of nitrogen, then on a hotplate at 70°C for 1 minute. Finally, the resist mask is cleaned in an oxygen plasma asher for 10 seconds, at a pressure of 0.2 hPa and an ionizing power of 75 W.
@@
For evaporating Josephson junctions, a good vacuum is required (at least below $5\cdot 10^{-7}$ hPa). Given that we are using the Manhattan process, the evaporator must be equipped with a planetary rotation. The parameters given here apply to one of our Plassys MEB 550S systems.
First, the surface is prepared by argon-ion milling with an incidence angle of 0°. The ion gun is tuned with the filament at 500 V and 25 mA, and the acceleration grid at 100 V. The shutter is removed from the ion-beam trajectory twice for 5 seconds each, with 1 minute of rest between exposures. Then comes the evaporation of the first layer of aluminum, performed at an angle of 30° and a rate of 1 nm/s. A thickness of 29 nm is set on the evaporator to obtain an actual thickness of 25 nm. This first layer is oxidized by a mixture of oxygen and argon at 10 hPa for 5 minutes. The sample holder is rotated around the planetary axis by 90° in order to evaporate along the orthogonal direction of the Manhattan pattern. Finally, the second evaporation of aluminum is performed with the same parameters (30° angle and 1 nm/s rate), except that a thickness of 69 nm is is set to get an actual thickness of 60 nm.
@@
Appendix D : wiring diagram
The wiring scheme used for SMPD characterization is presented in fig. 9. Lines ”In A1” and ”Out 2” are used to probe the buffer resonator in reflection (to determine $\omega_{b}$ and $\kappa_{b}$), and also to calibrate the detector efficiency by sending pulses containing a well-controlled number of photons. ”In B1” and ”In B4” are flux bias lines for controlling the SQUID inductances of the buffer and its Purcell resonator, respectively. ”In A2” and ”Out 3” are used for probing the waste resonator in reflection (to determine $\omega_{w}$ and $\kappa_{w}$), and also to perform dispersive readout of the qubit through the waste resonator. ”In B2” is used to pump the travelling-wave parametric amplifier (TWPA) on the ”Out 3” line. ”In B3” is the qubit drive line.
Figure 9: Wiring used for the characterization run reported herein.